OpenStudy (anonymous):

There is a single point which is a solution to the linear equation y-3=1/2(x+4) that is obvious because the equation is in point slope form. There are an infinite number of points that are not so obvious. Find the coordinates of the obvious point, and then find two additional points.

OpenStudy (anonymous):

y−3= 1 2 (x+4) y=0.5x+5

OpenStudy (anonymous):

@mathwiz916 wait, what??

OpenStudy (anonymous):

im back

OpenStudy (anonymous):

so..?

OpenStudy (anonymous):

@mathwiz916 I'm confused on your response.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

@mathwiz916 could you help me more? what are the points?

OpenStudy (anonymous):

google this y=0.5x+5

OpenStudy (anonymous):

it will show u the graph

OpenStudy (anonymous):

@mathwiz916 okay i did that.. but what does that have to do with what i'm asking??

OpenStudy (anonymous):

the points are -10 and 5

OpenStudy (anonymous):

So I would put (-10,5) and (4,-3) ? @mathwiz916 I need 2. Sorry if I sound stupid lol.

OpenStudy (anonymous):

here idk how to realy like explain so @welshfella can u help her out?

OpenStudy (anonymous):

@mathwiz916 sorry ;/ and thank you.

OpenStudy (anonymous):

lol dont be sorry its my fault

OpenStudy (welshfella):

its a funny question i would say that the point (-4,3) is the most obvious :- x = 4 when y = 3 because -4 makes the right side zero and 3 makes left side 0 3 - 3 = 0.5(-4+4) 0 = 0

OpenStudy (welshfella):

you can find 2 more points by putting x = 0 and finding y , and vice versa

OpenStudy (welshfella):

* typo x = -4 when y = 3

OpenStudy (anonymous):

@welshfella math is my worst enemy right now.. I got (-4,3) but I don't get how to get 2 additional points..

OpenStudy (welshfella):

y - 3 = 0.5(x + 4) let x = 0 y - 3 = 0.5(0 + 4) y - 3 = 0.5*4 y - 3 = 2 so what is y ?

OpenStudy (welshfella):

and to get another point plug in y = 0 into the equation and solve for x

OpenStudy (welshfella):

y - 3 = 2 add 3 to both sieds y - 3 +3 = 2 + 3 y = 5 so another point is (0,5)

OpenStudy (welshfella):

when y = 0 0 - 3 = 0.5(x + 4) -3 = 0.5x + 2 solve for x

OpenStudy (welshfella):

gotta go right now

OpenStudy (anonymous):

@welshfella omg! Thank you so much for your help!